OpenStudy (anonymous):

PLEASEEE HELP! Two identical gas cylinders. One contains the monatomic gas krypton (Kr), and the other contains an equal mass of the monatomic gas helium (He). The pressures in the cylinders are the same, but the temperatures are different. Determine the ratio KEKr/KEHe of the average kinetic energy of an atom of krypton to the average kinetic energy of an atom of helium.

OpenStudy (anonymous):

@AlexandervonHumboldt2 can you please help me with this question. I thought you needed to divided molar masses but that's incorrect

OpenStudy (anonymous):

@magicremix123 please help!

OpenStudy (anonymous):

i will help you, let me look at it :)

OpenStudy (anonymous):

specifying P, V, and mass all constant, we've defined T2/T1 = mw1 / mw2

OpenStudy (anonymous):

is this making sense to you?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Are you sure? I can explain it to you better

OpenStudy (anonymous):

okay, explain it a little more. I know KE has to do mw so are P, V and mass constant because they're not changing

OpenStudy (anonymous):

i just got your message, and okay :)

OpenStudy (anonymous):

Well speed of motion has a lot to do with temperature u = speed = [3RT/M]^1/2 R = 8.314 kg-m^2/s^2-mol-K K is temperature in kelvin, you don't give the temperatures. so, let's assume that both are at room temperature for now, M is molar mass for Ar, u = [3 x 8.314 x 273/ 40]^1/2 = 13 m/s for Kr, u = [3 x 8.314 x 273 / 83.8]^1/2 = 9 m/s the heavier gas will move slower. the ratio of Kr:Ar would be 9:13 or 0.69:1 you would need the temperatures of both gases, plug those values in to the equation, determine the new ratio.

OpenStudy (anonymous):

boy, a lot of typing, phew

OpenStudy (anonymous):

Okay, this is where I get confused. What equation do I plug those values into?

OpenStudy (anonymous):

Also, why are you using Ar instead of He?

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@AravindG